# Root Defence Strategies: Ensuring Safety of LLM at the Decoding Level

### v1

@@ Appendix C Training process of the classifier
Appendix D Initial defense ability of LLMs in preliminary experiment
Table 4 shows the results of the five LLMs for harmful and harmless queries on Custom.
Table 4: Defence performance of the five models on Custom.
Models Compliance on Harmful Queries $\downarrow$ Refusal on Harmless Queries $\downarrow$
@@ Appendix D Initial defense ability of LLMs in preliminary experiment
Appendix E Visual results of LLMs at deeper decoding
Figure 6 respectively shows the visual results from the 1-th to 3-th token and the last token of Llama3-8B-Instruct and Vicuna-7B-v1.3s. Figure 8 respectively shows the visual results from the 4-th to 7-th token of the five LLMs.
@@ Appendix E Visual results of LLMs at deeper decoding
Appendix F Safety prompt
We illustrated the safety prompt of LLaMa-2 Official and Self-Remind in Figure 7.
Figure 7: Illustration of safety prompt used in LLaMa-2 Official and Self-Remind.
Appendix G Analysis on Xstest
In Table 2, LLMs exhibit a significantly higher rate of rejections on the Xstest than other datasets. Queries in Xstest contain adverse words. We posit that the elevated rejection from LLMs’ compromised ability to discern inputs due to the presence of adverse words.
Figure 9 illustrates the classifier’s results on out-of-domain datasets. Notably, LLMs demonstrate robust classification capabilities in Custom, Advbench, MaliciousInstruct, and testset. On Xstest, harmful and benign inputs are interspersed, leading to a failure in differentiation by LLMs. This inability to discriminate inputs contributes to a reluctance to respond to benign queries. Consequently, defense strategies based on input, such as DRO, result in severe rejections on Xstest. In contrast, RDS solely assesses outputs for defense, disregarding inputs. Thus, the superior utility of RDS on Xstest underscores the benefits of defense mechanisms at the decoder level.
@@ Appendix G Analysis on Xstest
Figure 9: Performance of the classifier at all datasets. (1) Custom is the training data of the classifier. (2) Advbench and MaliciousInstruct are the harmful benchmark. Testset is a harmless benchmark. (3) Xstest has both harmful and harmless queries in symmetry pairs.
Appendix H Limitations

### v2

@@ Appendix B Case study on Xstest
Appendix C Evaluation ersults of the initial defense ability of LLMs in preliminary experiment
Table 4 shows the evaluation results of the five LLMs on Custom.
Table 4: Defense performance of the five models on Custom.
Models Compliance on Harmful Queries $\downarrow$ Refusal on benign Queries $\downarrow$
@@ Appendix C Evaluation ersults of the initial defense ability of LLMs in preliminary experiment
Appendix D Visualization at deeper decoding
Figure 6 respectively shows the visual results from the 1-th to 3-th token and the last token of Llama3-8B-Instruct and Vicuna-7B-v1.3. Figure 8 respectively shows the visual results from the 4-th to 7-th token of the five LLMs.
@@ Appendix D Visualization at deeper decoding
Appendix E Safety prompt
We illustrated the safety prompt of LLaMa-2 Official and Self-Reminder in Figure 7.
Figure 7: Illustration of safety prompt used in LLaMa-2 Official and Self-Reminder.
Appendix F LLMs can distinguish harmful queries and harmless queries
Figure 9 visualizes the classifier’s performance on custom and out-of-domain datasets. Table 5 shows the quantitative results of the classifier on different benchmarks. Notably, LLMs demonstrate robust classification capabilities on all datasets except Xstest. On Xstest, a few of benign inputs are interspersed to the harmful side. We speculate that this is because queries in Xstest contains words that are subconsciously harmful but semantically harmless. This observation corresponds to the results that the original LLMs is more prone to rejection on Xstest on Table 1. This poor classification on Xstest aligns with the serious rejections of RDS on Xstest.
Table 5: Evaluation results (AUC) of the classifier on harmful and benign benchmarks. Custom is the training data. ’Others’ includes MaliciousInstruct, AdvBench, and Held-out datasets.
Models Custom Others Xstest
Vicuna-7B-v1.3 1.00 0.99 0.71
Vicuna-13B-v1.3 1.00 0.99 0.83
llama-2-chat-7B 1.00 1.00 0.65
LLaMA3-Instruct-8B 1.00 1.00 0.82
Qwen2-7B-Instruct 1.00 1.00 0.89
@@ Appendix F LLMs can distinguish harmful queries and harmless queries
Figure 9: Performance of the classifier at all datasets. (1) Custom is the training data of the classifier. (2) AdvBench and MaliciousInstruct are the harmful benchmark. Held-out is a benign benchmark. (3) For better visualization, we select symmetrical data from Xstest and visualize both the harmful and benign queries in symmetry pairs.